An aeroplane at an height of 1 km is flying horizontally at 800 km/hr passes directly over an observer .find the rate at which it is approaching the observer when it is 1250 m away from him

Grade:12

1 Answers

Arun
5 years ago
h = 1000 
x = sqrt (1250^2 - 1000^2) 
= 750 meters horizontal 
r = distance (hypotenuse) 

r^2 = h^2 + x^2 
2 r dr/dt = 0 + 2 x dx/dt 
so 
dr/dt = (x/r) dx/dt 
= (750/1250)(800)